# Generator: A Long-Context Generative Genomic Foundation Model

### v1

@@ Abstract
Advancements in DNA sequencing technologies have significantly improved our ability to decode genomic sequences. However, the prediction and interpretation of these sequences remain challenging due to the intricate nature of genetic material. Large language models (LLMs) have introduced new opportunities for biological sequence analysis. Recent developments in genomic language models have underscored the potential of LLMs in deciphering DNA sequences. Nonetheless, existing models often face limitations in robustness and application scope, primarily due to constraints in model structure and training data scale. To address these limitations, we present Generator, a generative genomic foundation model featuring a context length of 98k base pairs (bp) and 1.2B parameters. Trained on an expansive dataset comprising 386B bp of eukaryotic DNA, the Generator demonstrates state-of-the-art performance across both established and newly proposed benchmarks. The model adheres to the central dogma of molecular biology, accurately generating protein-coding sequences that translate into proteins structurally analogous to known families. It also shows significant promise in sequence optimization, particularly through the prompt-responsive generation of promoter sequences with specific activity profiles. These capabilities position the Generator as a pivotal tool for genomic research and biotechnological advancement, enhancing our ability to interpret and predict complex biological systems and enabling precise genomic interventions.
@@ 1 Introduction
Figure 1: Overview of the Generator. (A) The pre-training dataset of the Generator encompasses a diverse range of eukaryotic organisms and gene types, totaling 386B nucleotides. (B) The pre-training employs the next token prediction (NTP) task, utilizing a 6-mer tokenizer. (C) Model comparison regarding the context length, pre-train data volume, and model parameters. (D) Downstream applications include sequence comprehension, interpretation of the central dogma by generating DNA sequences that encode functional proteins, and prompt-responsive promoter design with targeted activity profiles.
@@
Beyond benchmark performance, the Generator adheres to the central dogma of molecular biology [1], accurately generating protein-coding DNA sequences that produce proteins structurally analogous to known families. Moreover, the Generator showcases significant promise in sequence optimization, particularly in the design of promoter sequences that regulate gene activity during various biological stages, highlighting its potential for a series of biologically significant tasks. Our findings position the Generator as a vital resource for genomic research and biotechnological advancement. By enhancing our capability to interpret and predict genomic sequences, the Generator paves the way for profound improvements in our understanding of complex biological systems and the development of precise genomic interventions.
@@
    Demonstration of the model’s capability in promoter design, highlighting its potential to advance biotechnological applications and genomic research.
@@ 3.4.4 Sequence Design
In contrast, we employ a more efficient and flexible approach, refining supervised fine-tuning to accommodate prompt-responsive sequence generation. In our sequence design task, we utilize promoter sequences and activity data from the DeepSTARR [17] dataset. We select promoter sequences from the top and bottom quartiles of activity values. These samples are labeled with prompts <high> for high-activity and <low> for low-activity promoters. This refined fine-tuning approach enables the design of promoter sequences with desired activity profiles, illustrating the model’s capability in biologically directed sequence optimization.
@@ 4.3 Central Dogma
4.4 Promoter Design
We employed the promoter activity data from DeepSTARR [17], following the dataset split initially proposed by DeepSTARR and later adopted by NT. Using this data, we developed a promoter activity predictor by fine-tuning the Generator. This predictor surpasses the accuracy of DeepSTARR and NT-multi (Table 6), establishing itself as the current state-of-the-art predictor. By applying our refined SFT approach as outlined in Sec. 3.4.4, we generated a collection of candidate promoter sequences with specific activity profiles. As illustrated in Fig. 5, the predicted activities of these candidates exhibit significant differentiation between the generated high/low-activity promoters and natural samples.
@@ 4.4 Promoter Design
Figure 5: Promoter design. (A-B) Pearson correlation between the predicted promoter activity and the measured activity. (C-D) Distribution densities of the predicted activity of generated promoter sequences with distinct activity profiles.
@@ 5 Discussion & Future Development
A pivotal feature of the Generator is its adherence to the central dogma of molecular biology, generating protein-coding DNA sequences that can produce proteins analogous to those in established families, such as Histones and Cytochrome P450. This capability is validated by the structural congruence of translated proteins with existing structures, as verified by AlphaFold and Foldseek analyses. Additionally, in the sequence design task, the Generator demonstrates its capability in optimizing promoter sequences, achieving specific activity profiles with precision and computational efficiency. This underscores the model’s potential to advance the field of synthetic biology, facilitating precise sequence design and optimization and enhancing our understanding and manipulation of complex biological systems.
@@ Appendix D Details of Central Dogma
Appendix E Details of Promoter Design
For the promoter activity prediction, we adhered to the data partitioning of DeepSTARR [17]. We conducted the same hyperparameter search strategy as in our benchmark experiments and selected a learning rate of $2e^{-5}$ and a batch size of 64 for training the activity predictor. For promoter generation, we selected promoter sequences from the top and bottom quartiles of activity values within the DeepSTARR training set. These sequences were used to perform supervised fine-tuning (SFT) on the Generator, labeled with prompts <high> and <low> for high and low activity sequences, respectively. The hyperparameters for SFT were set to a learning rate of $1e^{-5}$ and a batch size of 2,048. During the generation process, we applied the same hyperparameter search strategy detailed in Sec. D. The hyperparameters used for visualization in Fig. 5 were set to $T=0.5$, $P=0.7$ for the developmental activity, and $T=0.8$, $P=0.6$ for the housekeeping activity.

### v2

@@ Abstract
Advancements in DNA sequencing technologies have significantly improved our ability to decode genomic sequences. However, the prediction and interpretation of these sequences remain challenging due to the intricate nature of genetic material. Large language models (LLMs) have introduced new opportunities for biological sequence analysis. Recent developments in genomic language models have underscored the potential of LLMs in deciphering DNA sequences. Nonetheless, existing models often face limitations in robustness and application scope, primarily due to constraints in model structure and training data scale. To address these limitations, we present Generator, a generative genomic foundation model featuring a context length of 98k base pairs (bp) and 1.2B parameters. Trained on an expansive dataset comprising 386B bp of eukaryotic DNA, the Generator demonstrates state-of-the-art performance across both established and newly proposed benchmarks. The model adheres to the central dogma of molecular biology, accurately generating protein-coding sequences that translate into proteins structurally analogous to known families. It also shows significant promise in sequence optimization, particularly through the prompt-responsive generation of enhancer sequences with specific activity profiles. These capabilities position the Generator as a pivotal tool for genomic research and biotechnological advancement, enhancing our ability to interpret and predict complex biological systems and enabling precise genomic interventions. Implementation details and supplementary resources are available at https://github.com/GenerTeam/GENERator.
@@ 1 Introduction
Figure 1: Overview of the Generator. (A) The pre-training dataset of the Generator encompasses a diverse range of eukaryotic organisms and gene types, totaling 386B nucleotides. (B) The pre-training employs the next token prediction (NTP) task, utilizing a 6-mer tokenizer. (C) Model comparison regarding the context length, pre-train data volume, and model parameters. (D) Downstream applications include sequence comprehension, interpretation of the central dogma by generating DNA sequences that encode functional proteins, and prompt-responsive enhancer design with targeted activity profiles.
@@
Beyond benchmark performance, the Generator adheres to the central dogma of molecular biology [1], accurately generating protein-coding DNA sequences that produce proteins structurally analogous to known families. Moreover, the Generator showcases significant promise in sequence optimization, particularly in the design of enhancer sequences that regulate gene expression during various biological stages, highlighting its potential for a series of biologically significant tasks. Our findings position the Generator as a vital resource for genomic research and biotechnological advancement. By enhancing our capability to interpret and predict genomic sequences, the Generator paves the way for profound improvements in our understanding of complex biological systems and the development of precise genomic interventions.
@@
    Demonstration of the model’s capability in enhancer design, highlighting its potential to advance biotechnological applications and genomic research.
@@ 3.4.4 Sequence Design
In contrast, we employ a more efficient and flexible approach, refining supervised fine-tuning to accommodate prompt-responsive sequence generation. In our sequence design task, we utilize enhancer sequences and activity data from the DeepSTARR [17] dataset. We select enhancer sequences from the top and bottom quartiles of activity values. These samples are labeled with prompts <high> for high-activity and <low> for low-activity enhancers. This refined fine-tuning approach enables the design of enhancer sequences with desired activity profiles, illustrating the model’s capability in biologically directed sequence optimization.
@@ 4.3 Central Dogma
4.4 Enhancer Design
We employed the enhancer activity data from DeepSTARR [17], following the dataset split initially proposed by DeepSTARR and later adopted by NT. Using this data, we developed an enhancer activity predictor by fine-tuning the Generator. This predictor surpasses the accuracy of DeepSTARR and NT-multi (Table 6), establishing itself as the current state-of-the-art predictor. By applying our refined SFT approach as outlined in Sec. 3.4.4, we generated a collection of candidate enhancer sequences with specific activity profiles. As illustrated in Fig. 5, the predicted activities of these candidates exhibit significant differentiation between the generated high/low-activity enhancers and natural samples.
@@ 4.4 Enhancer Design
Figure 5: Enhancer design. (A-B) Pearson correlation between the predicted enhancer activity and the measured activity. (C-D) Distribution densities of the predicted activity of generated enhancer sequences with distinct activity profiles.
@@ 5 Discussion & Future Development
A pivotal feature of the Generator is its adherence to the central dogma of molecular biology, generating protein-coding DNA sequences that can produce proteins analogous to those in established families, such as Histones and Cytochrome P450. This capability is validated by the structural congruence of translated proteins with existing structures, as verified by AlphaFold and Foldseek analyses. Additionally, in the sequence design task, the Generator demonstrates its capability in optimizing enhancer sequences, achieving specific activity profiles with precision and computational efficiency. This underscores the model’s potential to advance the field of synthetic biology, facilitating precise sequence design and optimization and enhancing our understanding and manipulation of complex biological systems.
@@ Appendix D Details of Central Dogma
Appendix E Details of Enhancer Design
For the enhancer activity prediction, we adhered to the data partitioning of DeepSTARR [17]. We conducted the same hyperparameter search strategy as in our benchmark experiments and selected a learning rate of $2e^{-5}$ and a batch size of 64 for training the activity predictor. For enhancer generation, we selected enhancer sequences from the top and bottom quartiles of activity values within the DeepSTARR training set. These sequences were used to perform supervised fine-tuning (SFT) on the Generator, labeled with prompts <high> and <low> for high and low activity sequences, respectively. The hyperparameters for SFT were set to a learning rate of $1e^{-5}$ and a batch size of 2,048. During the generation process, we applied the same hyperparameter search strategy detailed in Sec. D. The hyperparameters used for visualization in Fig. 5 were set to $T=0.5$, $P=0.7$ for the developmental activity, and $T=0.8$, $P=0.6$ for the housekeeping activity.